A geologist is camped 9,677 m (about 6 miles) from a volcano as it erupts.

(a) How much time elapses before the geologist hears the sound from the eruption? (Assume the air temperature is 20°C.)


(b) How much time does it take the seismic waves produced by the eruption to reach the geologist's camp, assuming the waves travel along the surface through granite as sound waves do? (Assume the granite temperature is 20°C.)
Answer #1

Given distance   d = 9677 m

speed of sound in air is v = 344 m/s

a)

the time elapsed t = d / v


t = 9677 / 344

t = 28.13 s

b)

the speed of sound in granite is given by v = 4000 m/s

therefore t = d / v

t = 9677 / 4000

t = 2.42 s